OpenStudy (anonymous):

Which ordered pairs in the form (x, y) are solutions to the equation (x + 6)y = 9? Choose all answers that are correct. A. (−5, 9) B. (−3, 3) C. (3, 1) D. (7, 9)

OpenStudy (anonymous):

i do not understand this plz help me i will give medal and fan!

OpenStudy (anonymous):

i want this question step by step so i know plz

jimthompson5910 (jim_thompson5910):

Let's check choice A (−5, 9) (x + 6)y = 9 (-5 + 6)*9 = 9 ... replace x with -5 and y with 9 (1)*9 = 9 9 = 9 ... this is true, so choice A (−5, 9) is a solution I'll let you check the others. There may be more than one answer

OpenStudy (anonymous):

thank you so much it may take me like 5 min i will tag u when i have figured everything out ;)

jimthompson5910 (jim_thompson5910):

alright

OpenStudy (anonymous):

i know b is wrong

OpenStudy (anonymous):

working on c and d

jimthompson5910 (jim_thompson5910):

why is B wrong? can I see your steps?

OpenStudy (anonymous):

c is correct

OpenStudy (anonymous):

ok one sec

OpenStudy (anonymous):

(3, 1) (x+6)y = 9 (3+6) * 1 = 9 (1) * 9 = 9

jimthompson5910 (jim_thompson5910):

B. (−3, 3)

OpenStudy (anonymous):

i did the same kind of steps for b and d and they were wrong is this right?

OpenStudy (anonymous):

(-3,3) (x+6)y = 9 (-3 + 6) * 9 = 27 (1) * 27 = 27 27 = 9 :( wrong

OpenStudy (anonymous):

oops, replace the 9 with one in the second eqution

jimthompson5910 (jim_thompson5910):

-3 + 6 is not equal to 1

OpenStudy (anonymous):

i mean 3 replace the 9 with 3

jimthompson5910 (jim_thompson5910):

and idk how you turned that 9 into 27

jimthompson5910 (jim_thompson5910):

and I'm not sure why you replaced y with 9 when y = 3

OpenStudy (anonymous):

OMG I DID 3 NOT -3!!!!!!!!!1

OpenStudy (anonymous):

thank you so much for catching that let me re work it is take one sec

OpenStudy (anonymous):

replace 9 with 27, which means b is right so it was a, b, and c

OpenStudy (anonymous):

(-3 + 6) * 3 = 9

jimthompson5910 (jim_thompson5910):

better

OpenStudy (anonymous):

sorry for the confusion, but is d is definetly wrong

jimthompson5910 (jim_thompson5910):

then you replace the `-3+6` part with `3` since `-3+6 = 3`

jimthompson5910 (jim_thompson5910):

D (7,9) (x + 6)y = 9 (7 + 6)*9 = 9 13*9 = 9 117 = 9 ... false so you are correct about D

OpenStudy (anonymous):

A B and C are correct i would not have known with ur help here is a medal!!!

jimthompson5910 (jim_thompson5910):

yes, points A,B,C all lie on the equation while D is some point that does not lie on the line

OpenStudy (anonymous):

do u agree @Shadowknight6263 ?

OpenStudy (anonymous):

yez